OpenStudy (anonymous):

What is the period of the function? y = 4 cos (2π x)

OpenStudy (anonymous):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

In general, we have this cosine equation y = A*cos(Bx - C) + D

jimthompson5910 (jim_thompson5910):

The period T is T = 2pi/B

OpenStudy (anonymous):

thanks for helping me, let me look at what you wrote hang on

OpenStudy (anonymous):

okay yes I remember learing that

jimthompson5910 (jim_thompson5910):

what is the value of B in this case?

OpenStudy (anonymous):

is it 4?

jimthompson5910 (jim_thompson5910):

B is the term in front of the x

jimthompson5910 (jim_thompson5910):

y = A*cos(Bx - C) + D

OpenStudy (anonymous):

oh then that's 1, b=1

OpenStudy (anonymous):

so 2pi/1 = 2pi

jimthompson5910 (jim_thompson5910):

y = A*cos(Bx - C) + D y = 4*cos(2pi*x - 0) + 0

jimthompson5910 (jim_thompson5910):

look again

OpenStudy (anonymous):

4 ?

jimthompson5910 (jim_thompson5910):

the 4 matches up with the A

jimthompson5910 (jim_thompson5910):

what matches up with the B?

OpenStudy (anonymous):

no I mean I tired to solve the equation, =4. but matches up with b? what do you mean? im sorry im a little confused

OpenStudy (anonymous):

2pi matches with b

jimthompson5910 (jim_thompson5910):

so B = 2pi

jimthompson5910 (jim_thompson5910):

T = 2pi/B T = 2pi/2pi ... plug in B = 2pi T = 1

OpenStudy (anonymous):

it sure does

jimthompson5910 (jim_thompson5910):

So the period is 1

OpenStudy (anonymous):

ahh I see

jimthompson5910 (jim_thompson5910):

In other words, the function repeats itself every 1 unit (say it has a peak at x = 2, it would have another peak at x = 3, then another at x = 4, etc etc)

OpenStudy (anonymous):

is that how I find all the periods? I mean with this sort of equation. the one you posted

jimthompson5910 (jim_thompson5910):

yes if you know the equation, you use a template like this

OpenStudy (anonymous):

can you help me with more qustions? or are you busy?

jimthompson5910 (jim_thompson5910):

it's slightly different with tangent functions though since it has a period of T = pi/B but it's the same with sine or cosines

jimthompson5910 (jim_thompson5910):

I can do one more

OpenStudy (anonymous):

which represents the reference angle for 5pi/6

jimthompson5910 (jim_thompson5910):

where is that angle? which quadrant?

OpenStudy (anonymous):

that's all the question says :/

jimthompson5910 (jim_thompson5910):

do you know how to convert that angle to degrees?

OpenStudy (anonymous):

not exactly, remind mee

jimthompson5910 (jim_thompson5910):

you multiply the radian angle by 180/pi

OpenStudy (anonymous):

esay enough

OpenStudy (anonymous):

so what is the radian angle...

OpenStudy (anonymous):

is it 6?

jimthompson5910 (jim_thompson5910):

5pi/6 is the entire radian angle

jimthompson5910 (jim_thompson5910):

(5pi/6)*(180/pi) = ???

OpenStudy (anonymous):

150 !

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

where is 150 degrees?

OpenStudy (anonymous):

is that in the angel? sounds like an angel to me

OpenStudy (anonymous):

I mean in a triangle

jimthompson5910 (jim_thompson5910):

150 degrees is an angle which quadrant is it in?

OpenStudy (anonymous):

im not quite sure I remember what a quadrant is

jimthompson5910 (jim_thompson5910):

http://www.mathsisfun.com/definitions/quadrant-graph-.html

jimthompson5910 (jim_thompson5910):

and here is the unit circle http://www.mathsisfun.com/geometry/unit-circle.html

OpenStudy (anonymous):

gahh okay let me take a looksie

OpenStudy (anonymous):

well its a positive number so a positive quadrent, how do I know if its x or y?

OpenStudy (anonymous):

quadrent 1

OpenStudy (anonymous):

forget my last question haha

jimthompson5910 (jim_thompson5910):

do you see where 0 degrees is? and 90 degrees?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

where is 90 degrees, mark it on this drawing 

jimthompson5910 (jim_thompson5910):

click on the pencil to draw on my drawing

OpenStudy (anonymous):